Circular No. 6793

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 J. Rhoads, Kitt Peak National Observatory (KPNO), for the KPNO γ-

ray-burst follow-up team, reports: “We imaged the field surrounding GRB 971214 with the KPNO 0.9-m telescope in the I band from Dec. 17.37 to 17.55 UT. The proposed optical counterpart (Halpern et al., IAUC 6788) was detected in a combined image with 220 min of total integration time. The transient was at mag I = 22.9 ± 0.4, based on stellar magnitudes and positions provided by J. Halpern (priv. comm.) for stars of I = 15.93 and 18.46 at α = 11^{h}56^{m}29^{s}.86, δ = +65^{o}12′15′′.9 (equinox 2000.0), and 11^{h}56^{m}25^{s}.80, +65^{o}11′35′′.9, respectively. Images and further details are

available at http://www.noao.edu/noao/grb/971214.html.”

4U 1915–05 D. Barret, J. F. Olive, and L. Boirin, Centre d’Etude Spatiale des Ray-

onnements, Toulouse; J. E. Grindlay, P. F. Bloser, and Y. Chou, Harvard-Smithsonian Center for Astrophysics; and J. H. Swank and A. P. Smale, Goddard Space Flight Center, NASA, report: “Observations of the x-ray burster 4U 1915–05, obtained with the RXTE PCA between 1996 Feb. and Oct., show clear evidence for high-frequency quasiperiodic oscillations (HFQPO) in its persistent x-ray emission. The strongest signals were detected as follows: May 16, 1016 ± 10 Hz, 6σ detection, FWHM = 75 ± 10 Hz, rms = 18 ± 3 percent, countrate 35.6 counts/s/PCU unit (5–30 keV); May 23, 830 ± 10 Hz, 7σ, 80 ± 10 Hz, 15 ± 2 percent, 34.8; June 1, 935 ± 10 Hz, 7σ, 70 ± 10 Hz, 15 ± 2 percent, 31.2; Sept. 6, 860 ± 20 Hz, 8σ, 200 ± 20 Hz, 17 ± 2 percent, 31.6. HFQPOs at comparable frequencies may also be present in the May 15 and May 17 data (890, 920 Hz), with lower significance. Furthermore, on May 16 we see evidence (3.3σ) for a second HFQPO peak at ∼ 650 ± 15 Hz (FWHM = 50 ± 20 Hz, rms = 12 ± 2 percent). A signal of 600 ± 10 Hz at 5.3σ is also detected in the Oct. 29 data (FWHM = 50 ± 10 Hz, rms = 11 ± 2 percent, 33.4 counts/s/PCU units) with no accompanying HFQPO detection above ∼ 900 Hz. For the two observations for which the countrates were significantly larger (March 13, 68.5 counts/s/PCU; May 5, 56.5), no HFQPOs were detected. Further analysis is continuing; nevertheless these results may be important for the planning of further RXTE observations of the source.”

1997 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green